Farmer McDonald gives banjo lessons for $20 an hour. One day, he spends 10 hours planting $100 worth of seeds on his farm. What opportunity cost has he incurred? What cost would his accountant measure? If these seeds yield $200 worth of crops, does McDonald earn an accounting profit? Does he earn an economic profit?


1
Expert's answer
2022-01-14T10:16:21-0500

"opportunity cost=10\times20=200"

explicit costs=100

explicit costs=200-100=100

 economic profit=200-100-200=-100

No economic profit